The initial number of atoms in a radioactive element is 6.0×1020 and its half-life is 10 hours. Calculated a) the number of atoms which have decayed in 30 hours. b) The amount of energy liberated if the energy liberated per atom decay is 4.0×10-13

Grade:12

2 Answers

Sandeep Singh Rawat
4 years ago
N=Noe-at,,,,,so number of atoms decayed is
No-N,,,,and a=0.693/t
Total energy liberated = energy liberated by one atom×no.of atom decayed
tej
2 years ago
N=Noe-at,,,,,so number of atoms decayed is
No-N,,,,and a=0.693/t
Total energy liberated = energy liberated by one atom×no.of atom decayed